OpenStudy (anonymous):

find an equation of the line containing points (-4,1) and (2,-3)

OpenStudy (anonymous):

First find the slope: (-3-1)/(2--4) => -4/6 = -2/3 Using y = mx + b substitute one point and the slope: 1 = -2/3(-4) + b solve for b 1 = 8/3 + b -5/3 = b y = -2/3 x - 5/3

OpenStudy (anonymous):

what you do is you use the slope formula y2-y1 ------ x2-x1 plug in the exponents like this: -3-1 ------ 2-(-4) (leave-change-change) =-4 /2 simplify: -2 (slope) m=-2 y=-2x=b substitite: -3=-2(2)+b -3=-4+b add'4' to get b alone four crosses out on the right 1=b so then go back to the equation and you get y=-2x+1

OpenStudy (anonymous):

Izzy 2-(-4) ≠ 2